# Re(l)ality:

### v3

@@ 5 Conclusion
This paper has employed the fiber bundle framework to examine the role of reference frames and relational observables in GR, highlighting the philosophical and practical consequences of two contrasting paradigms: the View from Nowhere and the View from Everywhere. Below, I summarise the key findings of the paper and their broader implications.
In Section 2, I used the fiber-bundle formalism to discuss reference frames and relational observables in GR. This approach defines the space of models as a principal bundle and introduces gauge orbits to describe symmetries within the theory. Reference frames, acting as section maps, facilitate the construction of gauge-invariant observables. This formalism elucidates the relationship between the choice of a reference frame and the choice of a gauge-fixing.

### v4

@@ 5 Conclusion
This paper has employed the fibre bundle framework to examine the role of reference frames and relational observables in GR, highlighting the philosophical and practical consequences of two contrasting paradigms: the View from Nowhere and the View from Everywhere. Below, I summarise the key findings of the paper and their broader implications.
In Section 2, I used the fibre-bundle formalism to discuss reference frames and relational observables in GR. This approach defines the space of models as a principal bundle and introduces gauge orbits to describe symmetries within the theory. Reference frames, acting as section maps, facilitate the construction of gauge-invariant observables. This formalism elucidates the relationship between the choice of a reference frame and the choice of a gauge-fixing.